# AuON: A Linear-time Alternative to Semi-Orthogonal Momentum Updates

### v2

@@ Abstract
Orthogonal gradient updates have emerged as a promising direction in optimization for machine learning.However, traditional approaches such as SVD/QR decomposition incur prohibitive computational costs of $\mathcal{O}(n^{3})$ and underperformed compared to well-tuned SGD with momentum, since momentum is applied only after strict orthogonalization. Recent advances, such as Muon, improve efficiency by applying momentum before orthogonalization and producing semi-orthogonal matrices via Newton–Schulz iterations, reducing complexity to $\mathcal{O}(n^{2})$. Nevertheless, quadratic costs remain a bottleneck.
In this work, we study the semi-orthogonal properties of momentum-based updates and develop a method to bound momentum updates under a spectral-norm trust region, preserving directional information without requiring explicit semi-orthogonalization.
We propose AuON (Alternative Unit-norm momentum updates by Normalized nonlinear scaling), a linear-time optimizer that achieves strong performance without constructing semi-orthogonal matrices, while preserving structural alignment and reconditioning ill-posed updates. Our approach combines hyperbolic-cosine RMS scaling transformations with normalization, demonstrating both effectiveness and computational efficiency compared to Newton–Schulz methods. We further introduce a hybrid variant (Hybrid-AuON) that applies a single Newton-Schulz iteration. Experiments across vision and language benchmarks show that AuON and its hybrid variant achieve performance comparable to strong baselines such as AdamW and Muon.
Code is available at: https://github.com/ryyzn9/AuON
@@ 1 Introduction
Optimization in deep neural networks remains a central challenge, particularly due to the ill-conditioning of gradient and momentum updates. Empirically, these updates often exhibit a high condition number, with most of the energy concentrated in a few dominant directions. In practical terms, the update vectors are nearly low-rank: a handful of directions dictate the optimization trajectory while many potentially informative directions may be suppressed. This imbalance reminds us of a squashed ball that can only roll efficiently along a single axis, ignoring other pathways that may be equally important for generalization and representation learning. One solution is to make all update directions unit length; recent work has proposed orthogonalization of gradients and momentum updates to achieve this property.
@@
By orthogonalizing an update matrix, we effectively discard the scaling information encoded in the singular values and modify each direction to enforce perpendicularity, redistributing the update length into unit vectors along different directions. In this sense, the resulting update behaves as a unit-norm update in the spectral domain, emphasizing the geometric structure of the optimization landscape rather than the raw gradient magnitudes. In simple terms, orthogonalization amplifies ’rare directions’ with small magnitude in the update but which are nevertheless important for learning. This perspective highlights how orthogonalization can prioritize exploration across all relevant directions, mitigating the dominance of a few high-energy components and facilitating more balanced learning dynamics Zhang et al. (2025). Orthogonalized updates can therefore be interpreted as spectral descent directions, ensuring that updates explore the space more evenly-crucial for generalization and representation learning.
@@
Figure 1: Visualization of the Newton–Schulz process (0.5) over 5 iterations, compared with AuON and Hybrid-AuON. The heatmaps (top) show progressive orthogonalization, with $MM^{\top}$ converging from a scattered structure (Step 0) to an identity-like diagonal (Step 5). The singular value plot (bottom) illustrates rapid convergence toward $1.0$, confirming orthogonalization.
@@
Tuddenham et al. (2022) proposed an approach for neural network optimization in which the gradient is first orthogonalized via singular value decomposition (SVD), followed by the application of momentum, and then the resulting momentum term is used as the update. They refer to this method as Orthogonal-SGDM. In their experiments, they observed that even in the best-performing configuration, Orthogonal-SGDM was outperformed by a well-tuned standard SGD with momentum. This is because applying momentum after strict orthogonalization damages the momentum mechanism: orthogonalizing gradients before accumulation prevents momentum from effectively reducing variance and maintaining beneficial directional information. Moreover, strict orthogonality erases singular-value magnitudes and over-constrains the step, collapsing its singular-value structure to an isometry. In effect, the update becomes a spectral-norm-constrained move that discards useful magnitude information, wiping out correlations between update directions. Making all updates unit length may also increase harmful alignment effects. Recent advances, such as Muon Jordan et al. (2024), improve efficiency and performance by producing a semi-orthogonal matrix using Newton–Schulz iterations rather than a full orthogonal matrix using SVD, and by reordering the momentum update to occur before semi-orthogonalization. This reduces complexity to $\mathcal{O}(n^{2})$, but quadratic costs still remain a bottleneck.
@@
In this paper, we focus on developing an alternative approach to bound updates with high condition numbers under a unit-norm constraint, without explicit Semi-Orthogonalization. Our goal is to achieve strong performance with $\mathcal{O}(n)$ time complexity, without compromising efficiency or speed. Empirically, we find that normalization followed by a hyperbolic-cosine scaling transformation yields promising results.
@@ 2.2 Semi-Orthogonalization
i.e., the Frobenius-nearest semi-orthogonal matrix that removes the amplitude information in $\Sigma$. Higham (2000)
@@
which is the linear minimization oracle (LMO) of a conditional-gradient step. Hence, the singular values are flattened; by contrast, on the standard spectral-norm ball, the LMO yields the rank-1 solution $u_{1}v_{1}^{\top}$.Lee (2021)
@@
In practice, $Q_{t}Q_{t}^{\top}$ is computed efficiently via a low-order Newton Schulz iteration, and $\alpha$ is chosen to match update RMS across shapes, enabling stability and learning-rate transfer. Semi-orthogonalization stabilizes training by bounding spectral energy and equalizing directional gains, preventing overshoot along sharp curvature, reducing oscillations, and enabling larger learning rates by decoupling orientation from scaleLiu et al. (2025)
@@ 2.3 Orthogonalized Momentum as a Spectral Trust-Region Method
Recent advances demonstrate that orthogonalized momentum in deep learning optimizers, particularly the Muon optimizer, admits a principled interpretation as the solution to a non-Euclidean trust-region subproblem under the spectral norm constraint Kovalev (2025). The core update rule can be formulated as
@@
Theoretical Advantages The orthogonalization-first approach provides superior variance reduction compared to alternative momentum–orthogonalization orderings. By applying orthogonalization to the momentum vector before the parameter update, the method preserves the accumulated directional information while eliminating scale-dependent instabilities. This design choice demonstrates both theoretical guarantees for convergence under non-convex objectives and empirical improvements in training stability across diverse architectures Liu et al. (2025).
@@ 3 Methods
We hypothesize that forcing all update directions to unit length can be problematic, as not all directions contribute equally to optimization progress some may be harmful (having a negative impact) or irrelevant to loss reduction. Our goal is to develop an alternative method that removes the harmful directions or alignments and preserves the beneficial properties of near semi-orthogonalization while selectively scaling directions under unit-norm: decrease the scales of rare update directions relative to dominant ones and keep them all under a unit-norm trust region, meaning prioritizing directions with favorable conditioning that correspond to well-conditioned subspaces of the loss landscape under a spectral-norm trust-region.One solution is to apply a temperature-scaled softmax update matrix, followed by L2-renormalization, to bound the step under a trust-region. But computing softmax may be problematic as it introduces computational bottlenecks, and it does not preserve the semi-orthogonal property that is needed for an optimizer like Muon . We empirically find that the normalization with hyperbolic functions (cosh ) helps us achieve spectral-norm trust-region Kovalev (2025) and helps us preserve the near semi-orthogonal like properties, that are more stable and equitable learning dynamics compared to conventional gradient-descent steps, and stabilizes training by bounding spectral energy into a unit vector and equalizing directional gains, preventing overshoot along sharp curvature, reducing oscillations, and enabling larger learning rates by decoupling orientation from scalePeletier and Schlichting (2023)
@@ 3.1 Nonlinear reshaping via hyperbolic cosine RMS scaling
Our main goal is to keep all the updated directions under unit spectral norm and remove the harmful directions. We empirically find out that the updated matrix divided by the scale factor of the RMS magnitude of cosh() helps us bound the dominant update directions that have a high condition number under unit spectral norm, and helps us to preserve the near semi-orthogonal-like properties. By doing this, we remove the harmful alignment and stabilize the updates and equitable learning dynamics. The overall equation is
@@
This encourages a spread of activations (diversity) without enforcing strict orthogo- nality Peletier and Schlichting (2023) and X is the updated momentum vector direction
@@
Define $\mathrm{rms}:=\|\cosh(\mathrm{update})\|_{F}/\sqrt{N}$, where $\cosh$ is applied only to compute a global, tail-sensitive scale Peletier and Schlichting (2022). Because $\cosh$ is even and rapidly increasing in $|x|$, heavy tails inflate $\mathrm{rms}$, which reduces the overall step size when forming $U:=\mathrm{update}/(\mathrm{rms}+10^{-8})$. Crucially, $\cosh$ is not applied to the propagated vector: $U$ is a uniform rescaling of $\mathrm{update}$, so the signs and all relative component ratios of $\mathrm{update}$ are preserved in $U$. This yields scale invariance with tail-aware damping, without introducing per-coordinate reweighting in the final update.
@@
  • Scale invariance. For any $c>0$, replacing $G$ by cGcG leaves $\mathrm{update}$ (and thus $U$) unchanged up to $\varepsilon$-terms.
@@ 3.2 Hybrid Approach
Figure 2: comparison of computation efficiency of different methods on (nxn) random matrices
@@
We therefore introduce a hybrid scheme: apply a single Newton-Schulz iteration to partially decorrelate the update, followed by AuON rescaling to enforce spectral contraction.
@@
  • Efficiency. Compared to Muon’s multi-step Newton-Schulz iterations (5-10 steps), Hybrid-AuON requires only one iteration, reducing overhead while retaining some correlation control.

### v3

# AuON: A Linear-time Alternative to Orthogonal Momentum Updates

@@ Abstract
Orthogonal momentum gradient updates have emerged to overcome the limitations of vector-based optimizers like Adam. The vector-based optimizer Adam suffers from high memory costs and ill-conditioned momentum gradient updates. However, traditional Orthogonal momentum approaches, such as SVD/QR decomposition, suffer from high computational and memory costs and underperform compared to well-tuned SGD with momentum.Recent advances, such as Muon, improve efficiency by applying momentum before orthogonalization and approximate orthogonal matrices via Newton-Schulz iterations, which gives better GPU utilization, active high TFLOPS, and reduces memory usage by up to 3x. Nevertheless, Muon(Vanilla) suffers from exploding attention logits and has cubic computation complexity. In this paper we , deep dive into orthogonal momentum gradient updates to find the main properties that help Muon to achieve remarkable performance.We propose AuON (Alternative Unit-norm momentum updates by Normalized nonlinear scaling), a linear-time optimizer that achieves strong performance without approximate orthogonal matrices, while preserving structural alignment and reconditioning ill-posed updates. AuON has an automatic (”emergency brake”) to handle exploding attention logits.. We further introduce a hybrid variant ( Hybrid-AuON)that applies the linear transformations with Newton-Schulz iterations which out performs Muon in the language modeling tasks. Code is available at: https://github.com/ryyzn9/AuON
@@ 1 Introduction
The pursuit of computational efficiency in training Large Language Models (LLMs) has catalyzed a shift away from purely scalar, coordinate-wise adaptive optimizers toward geometry-aware methods. For over a decade, the optimization landscape has been hegemonized by first-order adaptive moment estimation algorithms, principally Adam and its decoupled weight-decay variant, AdamW. These methods, while robust, operate under the simplifying assumption that the parameter space is Euclidean and that each parameter can be adapted independently based on its gradient history. However, as neural architectures have scaled into the regime of trillions of parameters—exemplified by models like the Mixture-of-Experts (MoE) limitations of this coordinate-wise paradigm have become increasingly stark. The inefficiencies manifest not merely in convergence speed but in the disconnect between the optimizer’s update steps and the underlying matrix structure of modern neural networks. Principally, Adam and its decoupled weight-decay variant, AdamW, suffer from ill-conditioning of gradient and momentum updates due to the momentum for a linear layer, naturally, a 2D matrix tends to become almost low-rank in practice.This means that only a small number of dominant directions really drive the updates. Though rare directions seem minor, they are often essential for effective learning and can help capture more nuanced patterns in the data. Empirically, these momentum updates often exhibit a high condition number, with most of the energy concentrated in a few dominant directions. In practical terms, the update vectors are nearly low-rank: a handful of directions dictate the optimization trajectory while many potentially informative directions may be suppressed. This imbalance reminds us of a squashed ball that can only roll efficiently along a single axis, ignoring other pathways that may be equally important for generalization and representation learning. One solution is to make all update directions unit length; recent work has proposed orthogonalization of gradients and momentum updates to achieve this property. Adam and its variant AdamW have another critical limitation: they require keeping two extra variables for every model parameter, meaning the optimizer state takes up about twice as much memory as the model itself. Furthermore, Adam treats all parameters as a single long vector, updating each value independently without considering any internal structure of the parameters.
@@
To overcome these issues, researchers proposed orthogonalizing the momentum gradient update matrix, which can effectively discard the scaling information encoded in the singular values and modify each direction to enforce perpendicularity, redistributing the update length into unit vectors along different directions. In this sense, the resulting update behaves as a unit-norm update in the spectral domain, emphasizing the geometric structure of the optimization landscape rather than the raw gradient magnitudes. In simple terms, orthogonalization amplifies ’rare directions’ with small magnitude in the update, but which are nevertheless important for learning. This perspective highlights how orthogonalization can prioritize exploration across all relevant directions, mitigating the dominance of a few high-energy components and facilitating more balanced learning dynamics zhang2025adagradmeetsmuonadaptive. Orthogonalized updates can therefore be interpreted as spectral descent directions, ensuring that updates explore the space more evenly-crucial for generalization and representation learning.
@@
Figure 1: Visualization of the Newton–Schulz process (0.5) over 5 iterations, compared with AuON and Hybrid-AuON(NS=1). The heatmaps (top) show progressive orthogonalization, with $MM^{\top}$ converging from a scattered structure (Step 0) to an identity-like diagonal (Step 5). The singular value plot (bottom) illustrates rapid convergence toward $1.0$, confirming orthogonalization.
@@
tuddenham2022orthogonalisinggradientsspeedneural proposed an approach for neural network optimization in which the gradient is first orthogonalized via singular value decomposition (SVD), followed by the application of momentum, and then the resulting momentum term is used as the update. They refer to this method as Orthogonal-SGDM. In their experiments, they observed that even in the best-performing configuration, Orthogonal-SGDM was outperformed by a well-tuned standard SGD with momentum. This is because applying momentum after orthogonalization which damages the momentum mechanism: orthogonalizing gradients before accumulation prevents momentum from effectively reducing variance and maintaining beneficial directional information. Moreover, orthogonality erases singular-value magnitudes and over-constrains the step, collapsing its singular-value structure to an isometry. In effect, the update becomes a spectral-norm-constrained move that discards useful magnitude information, wiping out correlations between update directions. Making all updates to unit length may also increase harmful alignment effects. Recent advances, such as Muon jordan2024muon, improve efficiency and performance by approximating an orthogonal matrix using Newton–Schulz iterations rather than a full orthogonal matrix using SVD, and by reordering the momentum update to occur before orthogonalization. This makes the Muon computational efficient in GPU, reduces the wall-clock time, and gives much higher throughput, though Muon has cubic complexity, is 2 x faster than Adam and AdamW, but still suffers from exploding attention logits. We find out that, by bounding the momentum updates under a spectral-norm trust region can preserve the directional information without requiring explicit orthogonalization
@@
In this paper, we focus on developing an alternative approach to bound updates with high condition numbers under a unit-norm constraint, without explicit Orthogonalization. Our goal is to achieve strong performance with without compromising efficiency or speed. Empirically, we find that normalization followed by a hyperbolic-cosine scaling transformation can bound the updates with high condition numbers under spectral-norm trust region and uses an automatic ”emergency brake” to significantly more resistant to exploding attention logits than Muon, especially in heavy‑tailed regimes.
@@ 2.2 Semi-Orthogonalization
i.e., the Frobenius-nearest semi-orthogonal matrix that removes the amplitude information in $\Sigma$. article2
@@
which is the linear minimization oracle (LMO) of a conditional-gradient step. Hence, the singular values are flattened; by contrast, on the standard spectral-norm ball, the LMO yields the rank-1 solution $u_{1}v_{1}^{\top}$.lee2021_vonneumann
@@
In practice, $Q_{t}Q_{t}^{\top}$ is computed efficiently via a low-order Newton Schulz iteration, and $\alpha$ is chosen to match update RMS across shapes, enabling stability and learning-rate transfer. Semi-orthogonalization stabilizes training by bounding spectral energy and equalizing directional gains, preventing overshoot along sharp curvature, reducing oscillations, and enabling larger learning rates by decoupling orientation from scaleliu2025muonscalablellmtraining
@@ 2.3 Orthogonalized Momentum as a Spectral Trust-Region Method
Recent advances demonstrate that orthogonalized momentum in deep learning optimizers, particularly the Muon optimizer, admits a principled interpretation as the solution to a non-Euclidean trust-region subproblem under the spectral norm constraint kovalev2025understandinggradientorthogonalizationdeep. The core update rule can be formulated as
@@
Theoretical Advantages The orthogonalization-first approach provides superior variance reduction compared to alternative momentum–orthogonalization orderings. By applying orthogonalization to the momentum vector before the parameter update, the method preserves the accumulated directional information while eliminating scale-dependent instabilities. This design choice demonstrates both theoretical guarantees for convergence under non-convex objectives and empirical improvements in training stability across diverse architectures liu2025muonscalablellmtraining.
@@ 3 Methods
We hypothesize that forcing all update directions to unit length can be problematic, as not all directions contribute equally to optimization progress some may be harmful (having a negative impact) or irrelevant to loss reduction.At first glance, it seems impossible: how can one suppress ”harmful” directions without knowing what they are in advance? If one never explicitly decorrelates or identifies bad modes , how does it preferentially suppress them?
The answer is that AuON exploits a statistical property of bad directions: they tend to be spiky.
Most directions contribute equitably to the loss landscape; harmful or noisy directions are statistically rare and often appear as outliers—components that are much larger than average. When a harmful direction emerges, it appears as a spike or heavy-tailed component in the update matrix
Harmful directions produce spiky gradients—they are discontinuous, scattered, and their components are extremely non-uniform. Normal, beneficial updates are smoother and have more balanced component magnitudes.Not all spiky updates are harmful and not all rare direction are spiky but spiky directions carry the potentially highest risk of maximum harm.Spiky updates are high-variance ”gambles.” If the spike hits a sensitive parameter, it causes massive damage (high loss).
Our goal is to develop a method that suppress the harmful directions(spiky-gradient updates) or alignments and preserves the beneficial properties of orthogonalization while selectively scaling directions under unit-norm: one solution is to decrease the scales of rare update directions relative to dominant ones and keep them all under a unit-norm trust region, meaning prioritizing directions with favorable conditioning that correspond to well-conditioned subspaces of the loss landscape under a spectral-norm trust-region.To do that we can apply a temperature-scaled softmax update matrix, followed by L2-renormalization, to bound the step under a trust-region. But computing softmax may be problematic as it introduces computational bottlenecks and it can also dampens useful directions as a side effect of dampening which is crucial for effective learning. We empirically find that the normalization with hyperbolic functions (cosh ) helps us bound the momentum updates under spectral-norm trust-region kovalev2025understandinggradientorthogonalizationdeep and it is a clever statistical mechanism (the cosh-RMS scaling) that automatically dampens harmful directions as a side effect of dampening all spiky components as a result we get ”emergency brake” property which can handle exploding attention logits than Muon, especially in heavy-tailed regimes.”Emergency Brake” is an emergent property of the hyperbolic cosine (cosh) normalization step.
Figure 2: comparison of computation efficiency of different methods on (nxn) random matrices, where Hybrid-Auon(NS=1)
Which makes the training are more stable and equitable learning dynamics compared to conventional gradient-descent steps, and stabilizes training by bounding spectral energy into a unit vector and equalizing directional gains, preventing overshoot along sharp curvature, reducing oscillations, and enabling larger learning rates by decoupling orientation from scalePeletier_2023
@@ 3.1 Nonlinear reshaping via hyperbolic cosine RMS scaling
Our main goal is to keep all the updated directions under unit spectral norm and remove the harmful directions. We empirically find out that the updated matrix divided by the scale factor of the RMS magnitude of cosh() helps us bound the dominant update directions that have a high condition number under unit spectral norm, and helps us to preserve the orthogonal-like properties. By doing this, we remove the harmful alignment and stabilize the updates and equitable learning dynamics. The overall equation is
@@
This encourages a spread of activations (diversity) without enforcing strict orthogo- nality Peletier_2023 and X is the updated momentum vector direction
@@
Define $\mathrm{rms}:=\|\cosh(\mathrm{update})\|_{F}/\sqrt{N}$, where $\cosh$ is applied only to compute a global, tail-sensitive scale article. Because $\cosh$ is even and rapidly increasing in $|x|$, heavy tails inflate $\mathrm{rms}$, which reduces the overall step size when forming $U:=\mathrm{update}/(\mathrm{rms}+10^{-8})$. Crucially, $\cosh$ is not applied to the propagated vector: $U$ is a uniform rescaling of $\mathrm{update}$, so the signs and all relative component ratios of $\mathrm{update}$ are preserved in $U$. This yields scale invariance with tail-aware damping, without introducing per-coordinate reweighting in the final update. As a resust as soon as attention logits start to explode (creating gradient spikes), AuON automatically suppresses the update to near-zero. It effectively freezes the unstable layer rather than pushing it further.it is tail-aware. It uses the cosh function to detect when the optimization landscape is becoming unstable (spiky gradients) and reacts by shrinking the step size
@@
  • Scale invariance. For any $c>0$, replacing $G$ by $cG$ leaves $\mathrm{update}$ (and thus $U$) unchanged up to $\varepsilon$-terms.
@@ 3.2 Hybrid Approach
We therefore introduce a hybrid scheme: apply a single Newton-Schulz iteration(Auon+) or 3(AuON+3) or 5(AuON+5) Newton-Schulz iterations to partially decorrelate the update, followed by AuON rescaling to enforce spectral contraction.
@@
  • Efficiency. Compared to Muon’s multi-step Newton-Schulz iterations (5-10 steps), AuON requires only one iteration, reducing overhead while retaining some correlation control.
@@
3.3 Temperature–Scaled AuON: Sensitivity to Tensor Size and Heavy Tails
In temperature-scaled AuON, the cosh input scale $\gamma$ (parameterized by the exponent $\alpha$) and the RMS exponent $\beta$ are the primary hyperparameters determining how early and how aggressively the brake engages on heavy-tailed updates.
$\tilde{G}=\frac{G}{\lVert G\rVert_{F}},$
then applies a temperature
[…]
to obtain
$g_{\text{scaled}}=\gamma\,\tilde{G}$.
The brake statistic is computed as
$r=\Bigl(\mathbb{E}\big[\cosh^{2}\!\big(g_{\text{scaled}}\big)\big]\Bigr)^{\beta/2},$
and the update takes the form
$\Delta\theta=-\,\mathrm{lr}\,\frac{G}{r+\varepsilon}.$
In this AuON variant, the temperature is set to $\gamma=N^{0.48}$, where $N$ is the number of elements in the update tensor, and the hyperbolic–cosine RMS statistic is raised to the power $3.5/2$. These two modifications make the brake extremely sensitive to both tensor width and the presence of heavy-tailed updates.
The update pipeline is:
$g\;\xrightarrow{\text{momentum}}\;g_{\text{hat}}=\frac{g}{\|g\|+\varepsilon}\;\xrightarrow{\text{temperature}}\;g_{\text{scaled}}=N^{\alpha}g_{\text{hat}},\qquad\alpha=0.48,$ (4)
$s=\mathbb{E}\!\left[\cosh^{2}\!\left(g_{\text{scaled}}\right)\right],\qquad r=s^{\,3.5/2},\qquad\Delta\theta=-\,\text{lr}\;\frac{g}{\,r+10^{-8}}.$ (5)
For small arguments, $\cosh(x)\approx 1+x^{2}/2$, giving $s\approx 1+\mathbb{E}[g_{\text{scaled}}^{2}],$ but since $g_{\text{scaled}}$ has variance proportional to $N^{2\alpha}$, even the quadratic term grows with tensor size. For rare large entries, $\cosh(x)\sim\tfrac{1}{2}e^{|x|}$, so a tiny number of outliers with $|x|\gg 1$ dominate $s$ exponentially. Raising $s$ to the exponent $3.5/2$ further super-linearizes this effect: if tail magnitude doubles, the brake factor $r$ grows by more than a factor of two, causing the step norm to collapse sharply.
The temperature parameter $\alpha$ controls cross-layer behavior. Smaller $\alpha$ slows the growth of $N^{\alpha}$, keeping $g_{\text{scaled}}$ mostly in the mild regime ($|x|\lesssim 1$) even in wide layers. This yields more uniform braking across tensor shapes and makes AuON resemble a soft, global trust-region scaling. Larger $\alpha$ amplifies $N^{\alpha}$, causing $r$ to explode in wide layers, which dramatically reduces their effective learning rate and strongly suppresses any bursty or sharp gradients.